CREXAuxiliaryMeasurements-Aug2020

From PREX Wiki
Jump to: navigation, search


CREX Main << August 2020 CREX Run Plan


Schedule

  • This plan is evolving as we move along
Day, Date Shift Tasks (Responsible Party) Comments/Status
Saturday Sept 12 Owl Production DONE
Day Production DONE
Swing (Zeke/Pitt) Production DONE
Sunday Sept 13 Owl (Adam/Tao) Production. (No scans done.) 39 min ABU!
Day (Bob/Eric) Final Q2 measurement. Optics [1] Acc mostly down; Bob took escorted access and switched to Counting Mode
Swing (Zeke/Pitt) Optics [2] Sanghwa for Analysis and remote assistance. Mostly completed optics.
Monday Sept 14 Owl (Adam/Amali) Inelastic spectrum. Poletip study [3] Chandan for Analysis and remote assistance. Finished optics, inelastic spectrum, poletip study
Day (Thia/Bob) 10am - noon Beam studies 1: Hall in Controlled Access (CA) - Install blind At PMT, Blind main Detector; noon - 4:15pm Beam available to ready Halls - we take data with blinded tubes. Devi required for tube blinding
Swing (Weibin/Eric, Devi on-call) 4:15pm - 8:15pm Beam studies 2: Hall in CA - restore US PMTs, Compton e-Det work; 8:15pm - 9:15pm Beam recovery and onward delivery. Production.
Tuesday Sept 15 Owl (Adam/Amali) Pedestal Scans, PITA scan. Production.
Day (Rom/Eric) SAD Safety Meeting 9-11am https://primetime.bluejeans.com/a2m/live-event/tkcbphay ; Production if beam available 7:45am South Linac down, cold box trip in CHL1, recovery 11.5 hours
Swing (Devi/Siyu) Beam recovery and on to production on Ca-48. Compton rates too high, running straight.
Wednesday Sept 16 Owl (Bishnu/Victoria) Production Running straight.
Day (Simona/Eric/B.Yale - standby) 8am start Moller measurement. Sign check. Followed by raster and Target study Cameron supporting the raster study remotely and Bob locally.
Swing (Devi/Tao) Complete raster and target study. Production No Compton.
Thursday Sept 17 Owl (Victoria/Siyu) Production. No Compton.
Day (Rom/Weibin) Production. Pedestal scans, PITA scan. No Compton.
Swing (Zeke/Adam) Production No Compton.
Friday Sept 18 Owl (Victoria/Siyu) 6am SHUTDOWN CREX1 is done! Long live CREX2!
Day (Rom/Weibin) -----

Optics and Q2

  1. Q2 measurement (Done)
    • Repeat of Q2 measurement, with improved pedestal noise control.
    • Detector counting mode.
    • S0 trigger
    • production raster and beam position
  2. Optics Data (Done)
    • Sieve IN, dp=+1, 0, (maybe -1). Raster/no raster thin C, raster H20
    • Priority is 1 dp change (both arms!) plus backout, and 10 runs. Watercell runs should be long.
    • Assume 3 hours for magnet changes and 2 hours for data taking. ~ 1 shift
    • See grid on optics data spreadsheet
  3. Energy Measurement (Done)
    • Doug did a parasitic energy measurement Sun, Sep 13 at 20:24 [4].

Backgrounds

  1. Blinded At PMTs (Done)
    • Longish access to setup and back out, but short data taking period (production conditions).
    • Reconfigure AT PMTs to use blinding cap, attach to US PMTs.
    • This will be done during the long beam studies break on Monday
    • software blinding on the AT PMTs must be disabled
    • Check pedestal noise in counting and in integrating mode. We need low noise is the blinded tube, and we can't add noise to our main detector!
    • Integrating measurement for asymmetry measurement
    • (Do we want counting mode measurement for rate in blinded At tube?)
    • (yield only normalizable using table-gain gain curves).
  2. Blinded Main detector? (Done)
    • Need 2 hours to blind main detectors. Then another hour after meaurement to unblind.
    • short-ish yield runs in integrating mode.
      • Software blinding must be disabled on the blinded MD tubes for this study
    • short counting mode run to get relative rate
    • Then return. Keep AT blinded piggybacked for asymmetry measurement.
  3. Inelastic spectrum scan (Done)
    • Needs a detailed run plan
    • Counting mode. Current 0.5uA or so? Turn quartz HV off.
    • Measure dp spectrum form dp=nominal to dp=-25%, in 5% steps. Use LHRS only.
    • should scale all spectrometer magnets (including septum)
  4. Pole tip re-scattering (Done)
    • Counting mode (VDC tracking). Counting mode on detector is ok (and helpful) but not needed.
    • Q3 turned off.
    • Run three septum setpoints: nominal, -5%, and +5%. Need ~1M S0 trigger events per run (and more is better).

Detectors

  1. Linearity Studies (PITA + current scans)
    • These are ongoing - should continue regularly. One more, on last day.
    • Current scans for detector and BCM pedestals. Skip the BPMs, I think.
  2. Linearity LED tests
    • This will be studied post run, during SAD
    • detector crew were left off the SAD access list - Access will be approved. Reapply for access.

Other

  1. Sign check, done
    • perform at next Moller polarimeter measurement
    • run Moller, Compton, and parity DAQ for quick PITA scan [5]
  2. Target/Raster study
    • Set up for normal production mode. Call in experts (Bob M in person, Cameron remote).
    • With each raster change we start with a brief C12 run for safety, then do a brief Ca48 run.
    • Raster Scan Run Plan

Moller polarimetry

  1. One more measurement, done [6], [7]

Compton polarimetry

  1. Spectrum run
  2. Laser polarization changes - aborted
  3. linearity studies During SAD